OpenStudy (anonymous):

Differentiate the following: f(x)=x(2x-5)^3

OpenStudy (anonymous):

let x=u and v= (2x-5)^3 dy/dx = vu' + uv'

OpenStudy (anonymous):

So, I know I have the product rule and the chain rule that need to be applied here.

OpenStudy (anonymous):

yeah so what's the problem if u know how to do it

OpenStudy (anonymous):

Right. So i did dy/dx = dy/du*du/dx

OpenStudy (anonymous):

Where u = 2x-5 and y= x*u^3

OpenStudy (anonymous):

So then I get the following:

OpenStudy (anonymous):

d/du(x*u^3) * d/dx(2x-5)

OpenStudy (anonymous):

And then I have to apply the product rule to x*u3, right?

OpenStudy (anonymous):

So that would be as follows: x*d/du(u^3)+d/dx(x)*(u^3)

OpenStudy (anonymous):

u're making this soo long...u should just apply the formula i wrote above in the first place

OpenStudy (anonymous):

I'm doing it how the professor showed us. I think she wants us to use substitution on the test. She's VERY specific.

OpenStudy (anonymous):

Let me try it your way, though.

OpenStudy (anonymous):

Give me a minute to see what I get?

OpenStudy (anonymous):

Please?

OpenStudy (anonymous):

your formula gives me (2x-5)^3+3x(2x-5)^2

OpenStudy (mathmale):

Substitution? I associate substitution with integration, the oppositie of differentiation. Arqum's differentiation formula (the product rule) is correct.

OpenStudy (anonymous):

yeah i know how professors can be... ok so im going to solve it for u u=x v=(2x-5)^3 u'=1 v'=6(2x-5)^2 dy/dx = vu' + uv' =(2x-5)^3 (1) + x(6)(2x-5)^2 so yeah your answer seems correct just take common and blingidy u have your answer

OpenStudy (anonymous):

So when I do it your way, I completely miss doing the derivative of 2x-3. When I do it the professor's way, it doesn't reduce.

OpenStudy (anonymous):

The derivative of 2x-5 is 2..

OpenStudy (mathmale):

You could, if you wanted to, factor (2x-5) out of this expression. Try it.

OpenStudy (anonymous):

yes when you find out the value of v' that is differentiation of (2x-5)^3 u have to differentiate 2x-5 too

OpenStudy (anonymous):

(2x-5)^2(2x-5)+6x which is (2x-5)^2(8x-5)

OpenStudy (anonymous):

I feel like I always forget something. Would you guys mind showing me the substitution method?

OpenStudy (anonymous):

I can't figure out why I don't get the same answer. I must be subbing wrong.

OpenStudy (anonymous):

I've been trying to do f(x)=x(2x-5)^3 where u=2x-5 and y=x*u^3

OpenStudy (anonymous):

I can't figure out why that doesn't work or where I go wrong. Yes, it's a longer route. I'm afraid this is the way she wants us to go on the test.

OpenStudy (anonymous):

Am I choosing the wrong values for my inner/outer function or is it because there are actually three functions?

OpenStudy (mathmale):

I think you'd be better off just differentiating the given function in the usual way. Then, if you wish, we could try doing it using a substitution. If you insist on using a substitution first: y=u*v y ' = u*v ' + v*u ' where u=x, u ' = 1, v = (2x+5)^3 and v ' = 3(2x+5)^2*2= ?? Substitute these four components into the derivative equation, above.

OpenStudy (anonymous):

I agree that Arqum's way was so much easier. I won't argue that!

OpenStudy (anonymous):

yeah...

OpenStudy (anonymous):

u're over complicating things here...just tell your professor to teach you guys the easy way not to confuse ya'll

OpenStudy (anonymous):

Haha. She's not the type of professor you say that to =)

OpenStudy (anonymous):

She's so confusing at times (and our textbook is worthless) that I went as far as to get the echapter for another textbook that covers this.

OpenStudy (anonymous):

LOL... i know your feel bro.. ( u r a bro eh?)

OpenStudy (anonymous):

As recommended to me by another professor in the math study hall.

OpenStudy (anonymous):

No, a chick =P It's cool, though.

OpenStudy (anonymous):

I appreciate everyone's help. Thank you.

OpenStudy (anonymous):

No problem...